OpenStudy (anonymous):

Show all work in dividing ((y^2-y-12)/(y^2-6y+8)) / ((y^2+2y-15)/(y^2-7y+10)). Use complete sentences to explain how to simplify this expression.

OpenStudy (turingtest):

\[\frac{\frac{y^2-y-12}{y^2-6y+8}}{\frac{y^2+2y-15}{y^2-7y+10}}\]???

OpenStudy (anonymous):

yup, do what your heart says.

OpenStudy (turingtest):

note that\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}\]and that all of the quadratics above can be factored in this way:\[ay^2+ay+b\to(y+c)(y+d)\]

OpenStudy (anonymous):

ohhhhhhhhhhh its

OpenStudy (turingtest):

so by the rule above we can rewrite our problem:\[\frac{\frac{y^2-y-12}{y^2-6y+8}}{\frac{y^2+2y-15}{y^2-7y+10}}=\frac{y^2-y-12}{y^2-6y+8}*\frac{y^2-7y+10}{y^2+2y-15}\]now factor each quadratic; I'll do one:\[y^2-y-12=(y-4)(y+3)\]factor each quadratic in that way, I bet some things will cancel...